wiki:UC20210706

Version 1 (modified by trac, 16 months ago) ( diff )

--

APO 3.5-m Users Committee Telecon, 07/06/2021


Attending: Nancy Chanover (NMSU), Bill Ketzeback (APO), Jamey Eriksen (APO), Mukremin Kilic (OU), Moire Prescott (NMSU), Misty Bentz (GSU), Kal Kadlec (UW), Anne Verbiscer (UVa), Eric Bellm (UW), Joanne Clark (Seattle), Ben Williams (UW), Kevin Schlaufman (JHU), Russet McMillan (APO), Denise Stephens (BYU)

User feedback and comments from institutional representatives

  • FGCU - no report
  • Georgia State - nothing to report
  • NMSU - nothing to report
  • BYU - nothing to report
  • Oklahoma - nothing to report
  • Wyoming - no report
  • Colorado - no report
  • UVa - nothing to report
  • JHU - nothing to report other than an issue with the observing schedule, which has the wrong instrument listed for one of the JHU programs - Russet will fix it before the next published schedule update
  • NAPG - no report
  • Seattle - nothing to report
  • Washington - nothing to report

Telescope and Instruments Report

Jamey's detailed report is included below, followed by additional information discussed during today's call.

3.5-m Telescope and Instruments Highlights, 5/28/2021 - 7/01/2021

0) Overview

As the State of New Mexico transitioned to different tier levels of COVID restrictions and prepared to remove capacity limits on business on 7/1/21, APO was at essentially full staff throughout the entire month of June.

Visitors have been at APO on and off throughout the month of June.

Weather for the month of June had been warming, and has now transitioned to monsoonal. The internet slow-down issue seems to have been fixed. If it is still affecting any users at your sites, please let us know.

1) Telescope

The telescope is operational, however, it is continuing to experience seasonal harmonic oscillations and audio ringing, particularly for the altitude axis. We are monitoring and adjusting the control parameters as needed. A tertiary rotation issue will be addressed by doing a full mapping during shutdown at the end of July.

2) Instruments

DIS: Scattered light is still quite good for both the red and blue. Red camera temperatures are elevated, which may necessitate taking darks during cals. The issue with the detent system has been addressed and the fix seems to be holding.

TripleSpec and Agile: cold, operational and performing nominally.

Echelle: The inter-order light has stabilized at a level that may cause problems for some science programs, particularly those that require a large number of blue ThAr lines for calibration, or programs with strong emission line targets. During shutdown we will try another partial warming intervention without cracking the vacuum, which should produce some improvement that we hope will hold until cooler weather arrives.

NICFPS: The instrument is performing nominally with the occasional need to reset the controller due to corrupted images. Observers may want to consider restricting rotation to keep the instrument near the upright position, where it has fewer problems.

ARCTIC: The instrument is performing nominally. Ion pump currents are elevated, indicating that a replacement will be needed soon. We are scheduling vacuum servicing the week before summer shutdown and it should be available about the same time as return to science at the end of shutdown.

3) ARCSAT Will not be available during the 3.5m shutdown: 7/19 – 8/3.


Additional telescope and instrument discussion

Not many updates beyond what was provided by Jamey above. Since New Mexico has removed all COVID restrictions for businesses there are no more staffing limits at APO. They have had lots of visitors over the past month. It has been fairly rainy for June! We are not aware of any more internet slowdowns but let us know if users experience anything. The instruments are all behaving; there was a hiccup with DIS last weekend but Bill fixed it. The echelle inter-order light has stabilized, albeit in a non-optimal state. As a reminder, ARCSAT will not be available during the 3.5m shutdown.


Recent improvements for users

Nancy reported on a few recent updates that will be of interest to users. First, the three new narrow-band filters for ARCTIC have arrived. As a reminder, these are 3 large format filters centered at H-alpha (6563 A), OIII (5007 A), and [SII] (6723 A), each with FWHM ~ 75 A. They have been catalogued and added to the filter inventory (http://filters.apo.nmsu.edu/index.php) and are now available for use. Second, there are two new TUI scripts available for users to try out. (1) a twilight flat script for ARCTIC. The user must determine the initial exposure time for each filter through a series of test flats, plug the times in, and the script automatically increments or decrements the exposure times (depending on whether using it for evening or morning twilight flats). It works very well with broadband filters, and reasonably well with others. (2) ARCES calibration script. This gives the user choices of which cals to take; users can uncheck the things they don't want. It also gives the option to allow the lamps to warm up. It utilizes the recommended exposure times given on the ARCES web page. Both scripts can be downloaded from the TUI scripts repository: https://www.apo.nmsu.edu/35m_operations/TUI-images/scripts/. We welcome feedback from users of any of these items!


Discussion of summer shutdown plans

It will be a short summer shutdown since the 3.5m primary mirror will not be realuminized this year. Tasks will include washing the optics, servicing ARCTIC, and general maintenance of various parts of the telescope that are due for routine maintenance. We will also be upgrading one of the UPS's on the intermediate level for the NA1 side of equipment (i.e. ARCES).


KOSMOS Update

The detector cabling is finished. The team installed the CCD and hooked it up. It was under vacuum this past weekend. The detector will be powered on this week so they can take some images in the lab.


ARCTIC Diffuser Rotator Update

The TEG team made some adjustments to the rotator. The parts being made by the machine shop are finished and will be picked up this week, then it will be assembled. The plan is to return the hardware to APO in early August for reinstallation.


Q3 3.5m scheduling

There is one OPEN slot before the shutdown: July 12B (dark), but ARCTIC is not available then. [Note added in proof: this is now spoken for.] There is some other OPEN and DD01 time later in schedule. If people need A halves, check the DD01 time. There are a few full nights remaining, mostly bright. Per usual, people interested in requesting this time should email Ben, Russet, Nancy, and their institutional scheduler.

There are two programs that requested training for non-students. This quarter we did not schedule a specific training program for those individuals but their observing schedules were scheduled in early August to overlap so they can shadow each other's observations and their afternoon training meetings will be combined.

Nancy polled the group concerning their thoughts about remote versus on-site training. With the site opening back up to visitors, we want to encourage groups to come to the observatory for in-person training since it is a unique experience that cannot be perfectly emulated in a virtual setting. That said, the Training and Education Working Group had articulated the need for a remote training option even before COVID, so we will be taking another look at the observatory policies regarding training. Several excellent points were raised during the discussion:

  • There may be some students or faculty who have mobility issues or other personal constraints that would make it difficult to travel to the site for the in-person training activities. There are others who may be immunocompromised, for whom interstate travel is still not safe due to COVID. We should ensure that there is an option for those individuals so that they are not excluded from training (and ultimately remote observing) opportunities.
  • Some doubted the need for hybrid training (with some members of a group on site with others participating remotely), while others thought they would take advantage of that option if it were available.
  • Some highlighted that there is a significant cost associated with bringing a class group to APO, thus remote training would still be an attractive option.
  • Some stated that they still plan to bring classes to APO.
  • The remote postdoc/staff training option is still appealing because in some cases those individuals are hired for short-term appointments. They need to be trained promptly so they can begin work on their projects right away rather than waiting several quarters by the time they can travel to the site and get trained.

The observatory management will be discussing this in the near future and will notify the group if there are any changes to the current training policy.


Q3 ARCSAT scheduling

We received zero requests for Q3 at the time of the proposal deadline; one request came in after that. The telescope will not be scheduled during the 3.5m summer shutdown, and we are reluctant to schedule ARCSAT for week of July 25 (immediately following the shutdown) since the Observing Specialists will be focused on shutdown recovery. As a reminder, Q3 goes to the end of September, so users are encouraged to request observations for the fall semester or quarter during Q3 (Q4 may end up being oversubscribed).


3.5m Science Symposium

Plans are shaping up nicely for the virtual science symposium that will be held on 26-28 July. We received approximately 55 abstracts for talks and/or posters, and the Science Organizing Committee will be putting the program together later this week. [Note added in proof: the web site is still under construction but the program for the oral sessions is now available: http://astronomy.nmsu.edu/aposcisymposium21/program.html.] Thanks to everyone who registered for the meeting, and to those who encouraged their students and colleagues to register. It's going to be great! Although the nominal registration deadline has passed, late registrations (including poster abstracts) will be accepted through Wednesday July 21.


Requests for updates on APO papers

Nancy made her annual request to have all users notify her of any papers (or theses/dissertations) published from Jan. 2020 - present that made use of data from the 3.5m and/or ARCSAT. The deadline to get this information to Nancy is August 1, 2021.


ACTION ITEMS

Open action items from previous meetings:

  • UC reps - warn your users of the following potential issues: (1) slow connection speeds to the observatory, (2) pointing errors after a slew due to oscillations in the altitude drive, (3) increased inter-order light on ARCES, (4) check NICFPS images and let the obs-spec know if they are getting a blank frame or jumbled signal so they can reset some of the electronics. Status: CLOSED.

New action items from this meeting:

  • UC reps: notify your users about the new ARCTIC filters and the two new scripts that are available to try out.
  • UC reps: ask your department members to send Nancy the citations for any 3.5m/ARCSAT-related papers published since January 2020. Deadline is August 1.

All Other Business

None. Meeting adjourned at 11:15 MDT.


Next meeting

The next meeting will be on Tuesday August 3 at 10:30 am MDT.


Note: See TracWiki for help on using the wiki.